Problem 4: Consider the refrigeration system shown below. The refrigerant is R-22, and the refrigerant mass flow rate is 0.2In the textbook used in class the specific enthalpy in the state is correct for 160bar
Answer #1

-Evaporator 2 Condensed 2 1-22 2-3 = Isentropic Compression Can. Pressure heat . Addition - (P2 - Po Addition (hy ha) (P= P4)= x=1, it will be in saturation varpow regioni: So its temperature it self will be loc of 4 6 = P =3.8 bar h=390 kJ/kg T= 100 h = ht: 220 koleg from chart, when h= 220 kr/kg at loe x=0.18 I 10 boi 3.8 bar 3.8bot 10 bar 10°c 50°c 10% 39o en 22-KI 4300.2 [430-225] = 42 kW Compressed Evaporatory Ennalle Condensat 42kw 34kW 8kW 9 COP> Boom 4.25 Coen From the available chait o